105 100 90 75 Total Utility 55 30 0 2 3 4 5 6 7 Quantity (pizzas/week)

  • What’s the Total utility consuming two pizzas per week?
  • What’s the Total utility consuming six pizzas per week?
  • What’s the Total utility consuming seven pizzas per week?
  • What’s the Marginal utility consuming two pizzas per week?
  • What’s the Marginal utility consuming six pizzas per week?
  • What’s the Marginal utility consuming seven pizzas per week
  • Should you eat that 7th pizza per week? Why?
  • What’s the optimal amount of pizzas to eat per week?
Answer #1

Total utility (2 pizzas per week)=55

Total utility (6 pizzas per week)=105

Total utility (7 pizzas per week)=100

Marginal utility(2 pizzas per week)=55-30=25

Marginal utility(6 pizzas per week)=105-100=5

Marginal utility(7 pizzas per week)=100-105=-5

You should not eat 7th pizza per week because Marginal utility will be negative at 7th pizza

the optimal amount of pizzas to eat per week=6 pizza at total utility is highest